OpenStudy (anonymous):

A pizza parlor has two different sizes of circular pizzas. The smaller one has a diameter of 12 inches and the larger one has a diameter of 20 inches. What is the ratio of their areas? I already got the area of each, 113 and 314, but I can't figure out the ratio.

OpenStudy (anonymous):

My choices are 9:25 3:5 \[2\sqrt{3}: 2\sqrt{5}\] \[\sqrt{6}: \sqrt{10}\]

OpenStudy (anonymous):

@Whitemonsterbunny17 any idea?

whitemonsterbunny17 (whitemonsterbunny17):

Ooooh, I didn't learn Geometry yet. But maybe @suckerofmath can help. c:

OpenStudy (anonymous):

Oh ok, well thanks anyway :)

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

circular pizzas so \[\pi r ^{2}\]

OpenStudy (anonymous):

But thats to find the area @Danyboii I already have that. I just need the ratio

OpenStudy (anonymous):

oh wait D: i think its smaller over larger

OpenStudy (anonymous):

i think since they've mentioned the smaller one first so you have to write the smaller first the n the largest.

OpenStudy (anonymous):

Give me 5-7 minutes

OpenStudy (anonymous):

Okay @Jesstho.-. @Danyboii that's what you normally do, put the first over the second \[\frac{ 113 }{ 314 }\] But you can't really simplify this so that's why I'm thinking its one of the answers with a square root.

OpenStudy (anonymous):

We know that the diameter of the smaller pizza is 12 inches, so the radius should be 6 inches right?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

and we know that the area of a circle is pi*r^2, right?

OpenStudy (anonymous):

Yeah and that'll give you 113

OpenStudy (anonymous):

So are of the smaller pizza should be pi*6^2 = pi*36=36*pi

OpenStudy (anonymous):

So it's area should be pi*10^2= 100*Pi

OpenStudy (anonymous):

Now ratio of the areas of two pizzas canceling out pis It becomes 35/100 Which is equal to 9/25 and that's out answer

OpenStudy (anonymous):

I hope you understood that

OpenStudy (anonymous):

Oh okay, yeah I got it. Thank you!

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

oh so didnt find the radius XD was going to explain u that earlier but u said u found the area XD oh well :3